OpenStudy (anonymous):

what is the greatest possible erroe of 11 1/8 oz?

OpenStudy (anonymous):

The easy trick is to take the number, make it 0, add a decimal then as many 0's as there are decimal numbers, then a 5 at the end. In the case of fractions, you see these go by quarter, so divide the quarter in half and you get 1/8, that's as far as you could vary from the number is half of whatever unit you are measuring. Check out the link for a more thorough explanation, look for the pattern in my answers, and do the last ones yourself to make sure you can do it again on the test later! 45.98 mg = 0.005 mg 12 1/4 in =1/8 in 54.4 cm =0.05 cm 1 3/4 c =1/8 cup 3 ft = 0.5ft 9 g =0.5g 12.3 L =0.05L 15.575 mm = 0.0005mm 24 1/2 yd = 512 m 10 1/8 oz. 3 1/16 i